# Towards a theory of natural directed paths

### v3

@@ 5. Application
  • $\operatorname{\overrightarrow{\operatorname{cosk}}}^{\Sigma}$ is the $\sigma$-labelled directed coskeleton of [5, Section 3.3] which is a tweak of the $\sigma$-labelled coskeleton functor of the category of precubical sets, the latter being badly behaved by [5, Proposition 3.15]: it contains too many cubes and some of them have to be identified;
  • $\operatorname{cosk}_{1}^{\widehat{\square}_{S},\overline{\Sigma}}$ is the $\sigma$-labelled coskeleton functor of the category of symmetric transverse sets which is well behaved by [6, Theorem 3.1.15].
@@
This implies that the parallel composition with synchronization for process algebra, as it is formalized in [6] using the labelled coskeleton functor of the category of symmetric transverse sets, has a category of cube chains which gives the correct homotopy type of tame natural $d$-paths.

### v4

@@ 5 Application
  • $\operatorname{\overrightarrow{\operatorname{cosk}}}^{\Sigma}$ is the $\sigma$-labelled directed coskeleton of (ccsprecub, , Section 3.3) which is a tweak of the $\sigma$-labelled coskeleton functor of the category of precubical sets, the latter being badly behaved by (ccsprecub, , Proposition 3.15): it contains too many cubes and some of them have to be identified;
  • $\operatorname{cosk}_{1}^{\widehat{\square}_{S},\overline{\Sigma}}$ is the $\sigma$-labelled coskeleton functor of the category of symmetric transverse sets which is well behaved by (symcub, , Theorem 3.1.15).
@@
This implies that the parallel composition with synchronization for process algebra, as it is formalized in symcub using the labelled coskeleton functor of the category of symmetric transverse sets, has a category of cube chains which gives the correct homotopy type of tame natural $d$-paths.